A car of mass 1200kg falls a vertical distance of 24m starting from rest. What is the workdone by the the force of gravity on the car? ( g = 10$\frac{s}^2$)
1
Expert's answer
2018-10-11T10:33:09-0400

The work done is

W=E_2-E_1=mgh_2-mgh_1,

where h_2=0 – final height of the car. Therefore,

W=-mgh_1=-1200·10·24=-288000 J.

Work done by the force of gravity is negative.